OpenStudy (anonymous):

Express the exact value of csc 60 degrees, with a rational denominator.

OpenStudy (callisto):

\[sin60 = \frac{\sqrt{3}}{2}\]\[csc60 = \frac{1}{sin60} = \frac{1}{\frac{\sqrt{3}}{2}} =...?\] *Note: all angles here are in degrees

OpenStudy (anonymous):

\[\csc60=\frac{1}{\sin60}=\frac{1}{\frac{\sqrt3}{2}}=\frac{2}{\sqrt3}=\frac{2\sqrt3}{3}\]Do you think that works?